Three-circle, red-on-white is one distinctive pattern painted on ceramic vessels of the Anasazi period found at the Wind Mountain archaeological site. At one excavation, a sample of 176 potsherds indicated that 62 were of the three-circle, red-on-white pattern. How many ceramic potsherds must be found and identified if we are to be 95% confident that the sample proportion of such potsherds is within 5.0% of the population proportion of three-circle, red-on-white patterns found at this excavation site? (Round your answer up to the nearest whole number.)
Answer #1

Solution,

Given that,

n = 176

x = 62

\hat p = x / n = 62 / 176 = 0.352

1 - \hat p = 1 - 0.352 = 0.648

margin of error = E = 5.0% = 0.05

At 95% confidence level

\alpha = 1 - 95%  

\alpha = 1 - 0.95 =0.05

\alpha/2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

Sample size = n = (Z\alpha / 2 / E )2 * \hat p * (1 - \hat p )

= (1.96 / 0.05)2 * 0.352 * 0.648

= 350.50

Sample size = n = 351